OpenStudy (anonymous):

Prove that \[{1-\tan^2x \over 1+ \tan^2x }= 1-2\sin^2x\]

OpenStudy (savvy):

LHS: \[(\cos^2x - \sin^2x)/(\cos^2x + \sin^2x)\] numerator is Cos2x and denominator is 1 now Cos2x is\[1 - 2\sin^2x\] = RHS henc proved...

OpenStudy (anonymous):

I don't understand. Where did the identities come from?

OpenStudy (savvy):

put tanx = Sinx/Cosx in first step.....

OpenStudy (lgbasallote):

\(\large 1 + \tan^2 x = \sec ^2 x\) so .... \(\LARGE \frac{1 - \tan^2 x}{\sec^2 x} = \frac{1}{\sec^2 x} - \frac{\tan^2 x}{\sec^2 x} = \cos^2 x - \sin^2 x\) = \(\large 1 - \sin^2 x - \sin^2 x = 1 - 2\sin ^2 x = RHS\)

OpenStudy (savvy):

now, do u knw that cos(A+B) = cosAcosB - SinASinb

OpenStudy (lgbasallote):

alternate solution :D

OpenStudy (lgbasallote):

i'll let @Savvy explain his solution..carry on ^_^

OpenStudy (lgbasallote):

@order do you understand my solution? :)

OpenStudy (anonymous):

Igbasallote, I understand yours, but not why you got -tan^2x/sec^2x to be - sin^2x

OpenStudy (savvy):

put tanx = sinx/cosx and secx = 1/cosx

OpenStudy (lgbasallote):

sorry for that.. \(\LARGE \tan ^2 x = \frac{\sin^2 x}{\cos ^2 x}\) \(\LARGE \sec^2 x = \frac{1}{\cos^2 x}\) \(\LARGE \frac{\tan ^2 x}{\sec^2 x} = \frac{(\frac{\sin ^2 x}{\cancel{\cos^2 x}})}{(\frac{1}{\cancel{\cos^2 x}})}\)

OpenStudy (anonymous):

Thank you!! :D